OpenStudy (anonymous):

Jin teaches tai chi at a local community center. He charges a one-time $12 enrollment fee and the cost of each class is shown below. Let x represent the number of classes and y represent the cost of each class. Cost of Classes (not including enrollment fee) Number of Classes (x) 1 2 3 4 Cost of Classes (y) $13.00 $22.00 $31.00 $40.00 Based on this information, what will it cost, in dollars, for Mark to take 7 classes?

OpenStudy (texaschic101):

you need to find the pattern in the numbers.... 22 - 13 = 9 31 - 22 = 9 40 - 31 = 9 you see the pattern....difference of 9 in between each one... so if 4 classes is 40, then 5 classes is (40 + 9) = 49. 6 classes = (49 + 9) = 58 I think you can figure out the 7th class....and don't forget to add the enrollment fee when figuring out the total cost.